OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {((-1, 11), (0, 8), (1, 7), (2, 5)} A.(-1, 11) B.(0, 8) C.(1, 7) D.(2, 5)

OpenStudy (mathmate):

I only see one pair which is NOT a solution.

OpenStudy (anonymous):

can you explain please im confused about this problem in general :/

OpenStudy (mathmate):

The ordered pairs mean (x,y). You feed x into the function and see if the y calculated by the function matches the y given in the ordered pair. If there is a match, the ordered pair is a solution to the function. For example: y = -2x + 9 ; Let's check the ordered pair (-1,11). x=-1, so substitute in the function, y=-2(-1)+9=2+9=11. Since 11 matches the value of y given, (-1,11) is a solution to the function. You can continue and reject the ones that do not match.